OpenStudy (celinegirl):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 A.) y - 6 = 1/2(x + 3) B.) y = 1/2x - 6 C.) y + 3 = 1/2(x - 6) D.) x - 2y = 12

Nnesha (nnesha):

\[\large\rm y-y_1 =m (x-x_1)\] point slope from where m is slope and passes through the point \[\rm (x_1,y_1)\]

OpenStudy (texaschic101):

you figure it out yet by what Nnesha said ?

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = 1/2 (6,-3)...x1 = 6 and y1 = -3 now sub those numbers into the formula and what do you get ?

OpenStudy (celinegirl):

@texaschic101 oops just saw their post! it would be A right?

OpenStudy (texaschic101):

no.... (6,-3)....x1 = 6 and y1 = -3 slope(m) = 1/2 try again

Nnesha (nnesha):

(x,y) order pair remember the alphabetical order x comes first then y

OpenStudy (texaschic101):

follow the formula... y - y1...so what is y1 ?

OpenStudy (texaschic101):

y1 = -3 (see how the points are labeled...(x1,y1) y - (-3) = m.....m is your slope...so what is m ?

OpenStudy (texaschic101):

it is just a matter of subbing...can you do that ?

Nnesha (nnesha):

i think she can.